OpenStudy (anonymous):

Help? 3.) Ellen works for a high-speed rail company that wants to develop a new rail line. Ellen’s project is to find a train that is the second fastest in the world. The Shinkansen Bullet Train in Japan is reported to go as fast as 320 kilometers per hour. The TGV train in France can reach speeds of 89.44 meters per second. Explain to Ellen how to find what speed her new train must go to be the second fastest in this group. Then, find an appropriate speed for Ellen’s train in miles per hour and use the formula distance = speed • time to find how far the train can go in 2.5 hours.

OpenStudy (anonymous):

You must find the spreed of BOTH trains in MILE PER HOUR. Explain to Ellen how to find what speed her new train must go to be the second fastest in this group. Then, find an appropriate speed for Ellen’s train in miles per hour and use the formula distance = speed • time to find how far the train can go in 2.5 hours

OpenStudy (tkhunny):

Well, do it. \(\dfrac{320\;km}{1\;hr}\cdot\dfrac{0.621371\;mile}{1\;km}\) That one is almost done. You finish this up and then do the other one.

OpenStudy (anonymous):

multiple that?

OpenStudy (tkhunny):

Look at the units. Most importantly, km/km leaves you with mile/hr. Are you suggesting you have never seen this, before?

OpenStudy (tkhunny):

0.621371 mi / 1 km = 1 -- It does not affect the value of the original expression. It just changes the units.

OpenStudy (anonymous):

I have all of the converted to km i'm just confused on how to do this part : Then, find an appropriate speed for Ellen’s train in miles per hour and use the formula distance = speed • time to find how far the train can go in 2.5 hours

OpenStudy (tkhunny):

What did you get for the speeds of the two trains in mph? And how did you get it?

OpenStudy (anonymous):

i have it converted to km not mph lol. I'm working on that part now .

OpenStudy (anonymous):

now i have it converted to miles

OpenStudy (tkhunny):

What did you get? 320 kilometers per hour = ?? mph 89.44 meters per sec = ?? mph

OpenStudy (anonymous):

320 Km = 198.839mph and 321.98 = 200.069mph

OpenStudy (anonymous):

well 89.44 = 200.069 miles

OpenStudy (tkhunny):

Very good. Now what?

OpenStudy (tkhunny):

So, 198.839 mph < Ellen < 200.069 mph ?

OpenStudy (anonymous):

okay, but what i'm confused on is they want me to fiind a speed for her and then tell them how far it goes in 2.5 hours

OpenStudy (tkhunny):

Agreed. That is a little confusing. I think it just needs to fall in that range. Just pick a speed. 200 mph? 199 mph? 199.438 mph? Just don't pick 198 mph or 201 mph. Multiply what you select by 2.5 and you're done.

OpenStudy (anonymous):

okay, I just hope I get it right tho i've had to go back multiple times and redo this and I can never understand it. >.< but thank you!

OpenStudy (tkhunny):

Nothing like a confusing question to make your life worse. Good luck.

OpenStudy (anonymous):

what was the answer